OpenStudy (anonymous):

Express answer in exact form. Find the area of the larger segment whose chord is 8" long in a circle with an 8" radius.

OpenStudy (anonymous):

195.26 sq/in

OpenStudy (zarkon):

\[\frac{160\pi}{3}+16\sqrt{3}\approx195.264421113\]